OpenStudy (anonymous):

Please help!!!!! A line passes through the point (-4, -6) and has a slope of 3. Write an equation for this line.

OpenStudy (anonymous):

y= 3 x +b you need to find b

OpenStudy (anonymous):

f(x)=3x + 6

OpenStudy (anonymous):

this is what I came up with: is the be 2? @eliassaab

OpenStudy (anonymous):

-6= 3(-4) + b b=6 y = 3x +1

OpenStudy (anonymous):

No, the 'b' is 6.

OpenStudy (anonymous):

wait doesn't 3*-4=-12

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

Yes, but you have to add -6 to both sides

OpenStudy (anonymous):

where in the world do you have add six to both sides?

OpenStudy (anonymous):

Then you have to subtract b to both sides which would get you -b, then you have to multiply each term by -1 to get rid of the negative sign on 'b'.

OpenStudy (anonymous):

you would divide 6 to both sides not add

OpenStudy (anonymous):

-6= 3(-4) + b -6 = -12 + b -6 + 6 = -12 + 6 + b 0 = -6 + b -b = -6 Multiply each term by -1: b = 6

OpenStudy (anonymous):

No, you would add, you only divide when you have a variable being multiplied to a number.

OpenStudy (anonymous):

y=3 x +6 It was a misprint in my previous post

OpenStudy (anonymous):

-6= 3(-4) + b b=6 y = 3x +6

OpenStudy (anonymous):

thanks guys

OpenStudy (anonymous):

No problem.